# Knowledge-Based Design Requirements for Generative Social Robots in Higher Education

### v1

@@ Abstract
Generative social robots (GSRs) powered by large language models enable adaptive, conversational tutoring but also introduce risks such as hallucinations, overreliance, and privacy violations. Existing frameworks for educational technologies and responsible AI primarily define desired behaviors, yet they rarely specify the knowledge prerequisites that enable generative systems to express these behaviors reliably. To address this gap, we adopt a knowledge-based design perspective and investigate what information tutoring-oriented GSRs require to function responsibly and effectively in higher education. Based on twelve semi-structured interviews with university students and lecturers, we identify twelve design requirements across three knowledge types: self-knowledge (assertive, conscientious and friendly personality with customizable role), user-knowledge (personalized information about student learning goals, learning progress, motivation type, emotional state and background), and context-knowledge (learning materials, educational strategies, course-related information, and physical learning environment). By identifying these knowledge requirements, this work provides a structured foundation for the design of tutoring GSRs and future evaluations, aligning generative system capabilities with pedagogical and ethical expectations.
@@ 1 Introduction
In the past, social robots were limited by pre-scripted dialogue and constrained interaction flexibility [30]. This made deployment challenging, given that students have diverse academic needs [56, 62]. More recently, human–robot interaction has advanced through generative AI, particularly Large Language Models (LLMs) [6, 10, 30]. Generative Social Robots (GSRs) use generative AI models such as LLMs to autonomously produce verbal and non-verbal communicative behavior [59]. This enables them to provide personalized and context-sensitive tutoring [1, 49, 50, 57, 56]. In doing so, GSRs may improve learning outcomes and student well-being by supporting motivation, reducing feelings of isolation during self-study, and providing accessible explanations and practice opportunities.
@@ 3 Methodology
The present research explores knowledge-based design requirements for tutoring-oriented GSRs in higher education, particularly to support interactive learning of lecture contents. Following Vonschallen et al. [60] – we conducted a qualitative interview study to identify desired GSR behaviors, as well as required self-, user-, and context-knowledge to interactively motivate and support students in learning lecture contents. The study was preregistered on AsPredicted.org (ID: #251,937) to enhance methodological transparency and limit post hoc changes in analytical focus and interpretation. The study was approved by the Ethics Review Board at Bielefeld University.
@@ 3.1 Sample
In total, twelve semi-structured interviews were conducted individually by the same member of the research team. Gender distribution of the interviewees was counterbalanced with six identifying as female, six as male. Eight out of the twelve participants reported prior experience in interactions with social robots. Four participants reported no prior interactions with social robots, however, two out of those four had observed such interactions. Lecturers had a mean age of 52 years (SD = 10.95 years). Most of them were rather experienced educators, with a minimum of six years of teaching, going up to 33 years (M = 18, SD = 8.94 years). The mean age of the student sample was 27 years (SD = 3.10 years). Student’s experiences ranged from the first semester of the bachelor’s degree (S5), up to the last semester in the master’s degree (S4). Table 1 provides an overview of the participants, as well as their teaching or study subjects.
@@ 3.2 Procedure
All interviews were conducted in German and lasted on average 45 minutes (SD = 12.65), ranging from 27 to 70 minutes. Interviews were audio-recorded and later transcribed verbatim for qualitative analysis.
@@ 4 Results
Participants mentioned a range of educational functionalities that the robot requires for effective use. All participants agreed that the robot should have motivational capabilities, specifically for persuading students to either start (S1, S4, S5) or to continue learning (S1, S5, L1, L4, L5). If students spend an unusually long time with uninterrupted learning, the robot should also suggest breaks, (S5, L6). Accordingly, the robot must have management skills, which should further be used to create learning plans (L4, L6, S3), and to give reminders (S1, S2, S4, S5, L4). Participants also uniformly stressed the need for adaptive learning support that aligns with the learning environment and is personalized to individual student needs. Importantly, the robot should provide personalized feedback, by tracking student goals and progress (S1, S2, S3, S4, L4) and by verifying information provided by the student (S4). The robot should also be able to create exercises (S1, S2, S3, S4) and illustrations (S1, S5) that fit the learning context. Further, it should clearly describe relevant procedures in applied courses, such as how to conduct a literature search (S2, L6).
@@
To responsibly conduct these tasks, participants mentioned a wide range of expected robot behaviors to ensure student safety. All participants emphasized the need for the robot to provide accurate information and to avoid hallucinations. Further, the robot should be transparent when retrieving information from the internet, and ask students for consent before doing so (S1, S2). To illustrate, S1 stated: “If it [the robot] realizes that it needs more context, then it may go on the internet – but it should state this and be transparent: I can’t get any further; is it okay if I do some research?” Transparency is also important in respect of student privacy (S2, L2, L3). When processing any student related data, students need to give informed consent and should have the option to deny access to private information (S2, S5, L2). To further guarantee user safety, the robot should avoid psychologically harmful actions (S1, S6, L1, L3, L4, L6) and biases regarding student gender, race, and intelligence (S1, L3, L4). S1 perceived this as an advantage of social robots compared to humans: “It is useful that students can ask [the robot] even the most basic questions without fear of being judged; this lowers the barrier for engagement”. When students mention personal topics such as fear of exams or stress, the robot should listen, rather than providing psychological advice (L6). Instead – if necessary – it should recommend psychological help (L4, L6). Participant L6, with a background in psychology explained it like this: “You shouldn’t go too deep into psychological topics, but rather listen and then move to the level of action. What can the student do with regards to their educational goals?” The need for a goal-oriented behavior that focuses on learning and avoids emotional attachment was also mentioned by other participants (S2, S3, L1, L3, L5). Particularly, the robot should not replace study groups, but rather be complementary (S1, L4, L5, L6). Most students and lecturers also highlighted a need for emotional robot engagement to provide learning motivation (S1, S4, S5, S6, L1, L2, L4, L5, L6). However, this motivating behavior should have clear boundaries regarding student autonomy (S1, S2, S3, L6). The robot should not be used to control students (S1) and students should be able to interrupt and shut down the robot at any time (S1, S2, S3). All students and some lecturers shared the view that the robot should be used primarily by students, not by lecturers (S1, S2, S3, S4, S5, S6, L2, L6). As such, the robot should not be used by lecturers to give mandatory tasks (S2, S3, S4, S5, S6, L6), to control students (S1), or to facilitate communication with students (S1).
@@ 5 Discussion
Regarding self-knowledge, our results highlight that a customizable role and personality is an essential prerequisite for a GSR to enable responsible and effective tutoring support in higher education. While most participants preferred a role of a study buddy, the preferred strictness of the robot varied from student to student. In particular, the robot should express balanced assertiveness that adapts to student needs. This aligns with prior research on the importance of assertive robot behavior in motivational contexts [3, 37, 40]. Friendliness – expressed through a patient, empathetic personality – was deemed essential for maintaining a positive learning environment, consistent with findings from educational HRI research [52]. Participants further agreed that the robot’s conscientiousness is key to provide competent, transparent and accurate learning support. This is supported by prior research demonstrating that robot competence is important to foster trust and to avoid pedagogical harm [12], especially in light of known risks of hallucinations in generative systems [13].

### v3

# arXiv:2602.12873v3[cs.HC] 02 Mar 2026

@@ Abstract
Generative social robots (GSRs) powered by large language models enable adaptive, conversational tutoring but also introduce risks such as hallucinations, overreliance, and privacy violations. Existing frameworks for educational technologies and responsible AI primarily define desired behaviors, yet they rarely specify the knowledge prerequisites that enable generative systems to express these behaviors reliably. To address this gap, we adopt a knowledge-based design perspective and investigate what information tutoring-oriented GSRs require to function responsibly and effectively in higher education. Based on twelve semi-structured interviews with university students and lecturers, we identify twelve design requirements across three knowledge types: self-knowledge (assertive, conscientious, and friendly personality with customizable role), user-knowledge (personalized information about student learning goals, learning progress, motivation type, emotional state, and background), and context-knowledge (learning materials, educational strategies, course-related information, and physical learning environment). By identifying these knowledge requirements, this work provides a structured foundation for the design of tutoring GSRs and future evaluations, aligning generative system capabilities with pedagogical and ethical expectations.
@@ 1 Introduction
In the past, social robots were limited by pre-scripted dialogue and constrained interaction flexibility [31]. This made deployment in educational settings challenging, given that students have diverse academic needs [57, 63]. More recently, human–robot interaction has advanced through generative AI, particularly Large Language Models (LLMs) [7, 11, 31]. Generative Social Robots (GSRs) use generative AI models such as LLMs to autonomously produce verbal and non-verbal communicative behavior [60]. This enables them to provide personalized and context-sensitive tutoring [1, 50, 51, 58, 57]. In doing so, GSRs may improve learning outcomes and student well-being by supporting motivation, reducing feelings of isolation during self-study, and providing accessible explanations and practice opportunities.
@@ 3 Methodology
The present research explores knowledge-based design requirements for tutoring-oriented GSRs in higher education, particularly to support interactive learning of lecture contents. Following Vonschallen et al. [61], we conducted a qualitative interview study to identify desired GSR behaviors, as well as required self-, user-, and context-knowledge to interactively motivate and support students in learning lecture contents. The study was preregistered on AsPredicted.org (ID: #251,937) to enhance methodological transparency and limit post hoc changes in analytical focus and interpretation. The study was approved by the Ethics Review Board at Bielefeld University.
@@ 3.1 Sample
A total of twelve semi-structured interviews were conducted between November 2024 and January 2025. This sample size is considered sufficient to achieve data saturation for qualitative analysis [2]. Gender distribution of the interviewees was counterbalanced with six identifying as female, six as male. Eight out of the twelve participants reported prior experience in interactions with social robots. Four participants reported no prior interactions with social robots, however, two out of those four had observed such interactions. Lecturers had a mean age of 52 years (SD = 10.95 years). Most of them were rather experienced educators, with a minimum of six years of teaching, going up to 33 years (M = 18, SD = 8.94 years). The mean age of the student sample was 27 years (SD = 3.10 years). Student’s experiences ranged from the first semester of the bachelor’s degree (S5), up to the last semester in the master’s degree (S4). Table 1 provides an overview of the participants, as well as their teaching or study subjects.
@@ 3.2 Procedure
All interviews were conducted individually by the same member of the research team in German and lasted on average 45 minutes (SD = 12.65), ranging from 27 to 70 minutes. Interviews were audio-recorded and later transcribed verbatim for qualitative analysis.
@@ 4 Results
Participants mentioned a range of educational functionalities that the robot requires for effective use. All participants agreed that the robot should have motivational capabilities, specifically for persuading students to either start (S1, S4, S5) or to continue learning (S1, S5, L1, L4, L5). If students spend an unusually long time with uninterrupted learning, the robot should also suggest breaks, (S5, L6). Accordingly, the robot must have time and task management skills, which should further be used to create learning plans (L4, L6, S3), and to give reminders (S1, S2, S4, S5, L4). Participants also uniformly stressed the need for adaptive learning support that aligns with the learning environment and is personalized to individual student needs. Importantly, the robot should provide personalized feedback, by tracking student goals and progress (S1, S2, S3, S4, L4) and by verifying information provided by the student (S4). The robot should also be able to create exercises (S1, S2, S3, S4) and illustrations (S1, S5) that fit the learning context. Further, it should clearly describe relevant procedures in applied courses, such as how to conduct a literature search (S2, L6).
@@
To responsibly conduct these tasks, participants mentioned a wide range of expected robot behaviors to ensure student safety. All participants emphasized the need for the robot to provide accurate information and to avoid hallucinations. Further, the robot should be transparent when retrieving information from the internet, and ask students for consent before doing so (S1, S2). To illustrate, S1 stated: “If it [the robot] realizes that it needs more context, then it may go on the internet – but it should state this and be transparent: I can’t get any further; is it okay if I do some research?” Transparency is also important in respect of student privacy (S2, L2, L3). When processing any student related data, students need to give informed consent and should have the option to deny access to private information (S2, S5, L2). To further guarantee user safety, the robot should avoid psychologically harmful actions (S1, S6, L1, L3, L4, L6) and biases regarding student gender, race, and intelligence (S1, L3, L4). S1 perceived this as an advantage of social robots compared to humans: “It is useful that students can ask [the robot] even the most basic questions without fear of being judged; this lowers the barrier for engagement”. When students mention personal topics such as fear of exams or stress, the robot should listen, rather than providing psychological advice (L6). Instead – if necessary – it should recommend professional psychological help (L4, L6). Participant L6, with a background in psychology explained it like this: “You shouldn’t go too deep into psychological topics, but rather listen and then move to the level of action. What can the student do with regards to their educational goals?” The need for a goal-oriented behavior that focuses on learning and avoids emotional attachment was also mentioned by other participants (S2, S3, L1, L3, L5). Particularly, the robot should not replace study groups, but rather be complementary (S1, L4, L5, L6). Most students and lecturers also highlighted a need for emotional robot engagement to provide learning motivation (S1, S4, S5, S6, L1, L2, L4, L5, L6). However, this motivating behavior should have clear boundaries regarding student autonomy (S1, S2, S3, L6). The robot should not be used to control students (S1) and students should be able to interrupt and shut down the robot at any time (S1, S2, S3). All students and some lecturers shared the view that the robot should be used primarily by students, not by lecturers (S1, S2, S3, S4, S5, S6, L2, L6). As such, the robot should not be used by lecturers to give mandatory tasks (S2, S3, S4, S5, S6, L6), to control students (S1), or to facilitate communication with students (S1).
@@ 5 Discussion
Regarding self-knowledge, our results highlight that a customizable role and personality is an essential prerequisite for a GSR to enable responsible and effective tutoring support in higher education. While most participants preferred a role of a study buddy, the preferred strictness of the robot varied from student to student. In particular, the robot should express balanced assertiveness that adapts to student needs. This aligns with prior research on the importance of assertive robot behavior in motivational contexts [4, 38, 41]. Friendliness – expressed through a patient, empathetic personality – was deemed essential for maintaining a positive learning environment, consistent with findings from educational HRI research [53]. Participants further agreed that the robot’s conscientiousness is key to provide competent, transparent, and accurate learning support. This is supported by prior research demonstrating that robot competence is important to foster trust and to avoid pedagogical harm [13], especially in light of known risks of hallucinations in generative systems [14].